OpenStudy (anonymous):

I'll become a fan and give a medal. 16^5/4 in simplest form

rishavraj (rishavraj):

use 16 = 2^4

OpenStudy (zzr0ck3r):

\(16^{\frac{5}{4}}=(16^\frac{1}{4})^5=(\sqrt[4]{16})^5=2^5= \ ?\)